OpenStudy (anonymous):

Let f (x) = x^3+x and let g=f^(-1) (you may take it as given that an inverse exists). Find g'(0) and g'(10).

OpenStudy (anonymous):

that was wrong \[\frac{d}{dx}[f^{-1}(x)]=\frac{1}{f'(f^{-1}(x))}\]

OpenStudy (anonymous):

you want the derivative of the inverse ( here called \(g\) ) at \(x=0\)so we need two things, \(f^{-1}(0)\) and \(f'(x)\)

OpenStudy (anonymous):

\[f^{-1}(0)\] solve \[f(x)=0\] for \(x\) via \[x^3+x=0\implies x=0\] so \[f^{-1}(0)=0\]

OpenStudy (anonymous):

\[f'(x)=3x^2+1\] and therefore \[f'(f^{-1}(0))=f'(0)=1\] and so your answer is \(\frac{1}{1}=1\)

OpenStudy (anonymous):

Thank you very much! Does that mean: To find g'(10): \[solve, x^3+x=10\] \[\therefore x=2\] \[f^{-1}(2)=13\] \[\therefore g'(10)=1/13\]

OpenStudy (anonymous):

x=y^3+y ,(x=0,y=0), (x=10,y=2) so 1=3y'y^2+y' then y'=1/(3y^2+1)= 1/1 , 1/13

OpenStudy (anonymous):

yes, you got it!